# Choosing alpha post hoc: the danger of multiple standard significance thresholds

### v1

@@ 4 Example 2: Infinitely many significance levels $\alpha$
Now suppose there are not two thresholds to choose from but many. To take this to the extreme, suppose each $0<\alpha\leq C$ can be chosen. Here $0<C\leq 1$ may for example be $0.1$. It is then easily imaginable that the author is biased towards choosing an $\alpha$ that is at least as large as her p-value $p$, when possible. To fix ideas, suppose the researcher in fact selects $\alpha=p$, unless $p>C$, in which case she selects $\alpha=C$. As usual, we denote the test by $\phi_{\alpha}$, where $\alpha$ depends on $p$.
In case $p\leq C$, we always have $p\leq\alpha$, so that $H_{0}$ is always rejected. Likewise, if $\alpha<C$, $H_{0}$ is always rejected. Thus, under $H_{0}$, conditional on $\alpha<C$, the type I error rate is not below $\alpha$. Specifically, for every $a<C$, we have $r_{a}=\mathbb{E}(\phi_{\alpha}|\alpha=a)/a=1/a$. Thus, the smaller $a$ is, the larger $r_{a}$ is, i.e. the larger discrepancy ratio is conditional on $\alpha=a$.
@@
Next, we may ask whether we have type I error control not conditionally, but at least on average, in the sense that $\mathbb{E}(r_{\alpha})\leq 1$ under $H_{0}$. Since $p$ is uniform on $[0,1]$ and $p\leq\alpha$ if and only if $p\leq C$, we have
$\mathbb{E}(r_{\alpha})=\mathbb{E}(\phi_{\alpha}/\alpha)=$
$\mathbb{P}(p>\alpha)\cdot 0+\mathbb{P}(p\leq\alpha)\mathbb{E}(\phi_{\alpha}/\alpha|p\leq\alpha)=$
$\mathbb{P}(p\leq\alpha)\mathbb{E}(\phi_{\alpha}/\alpha|\alpha=p\leq C)=$
$\mathbb{P}(p\leq\alpha)\mathbb{E}(1/p|p\leq C)=$
@@
We see that the expected discrepancy ratio $\mathbb{E}(r_{\alpha})$ is not only larger than 1, but in fact infinity. Thus, in expectation, the test is completely out of control.
@@ 5 Connection to e-values
A partial solution to the problem of data-dependent $\alpha$ are so-called e-values. An e-value is a measure of evidence that has been recently proposed as an alternative to the p-value (Howard et al., 2021; Shafer, 2021; Vovk and Wang, 2021; Grünwald et al., 2024; Ramdas et al., 2023; Koning, 2024b). An e-value is typically defined as a non-negative random variable $e$, whose expectation is bounded by 1 under the hypothesis $H_{0}$:
$\mathbb{E}\,e\leq 1.$ (5)
In the context of a simple null and alternative hypothesis that only contain a single distribution, a prominent example of an $e$-value is the likelihood ratio between these distributions (Grünwald et al., 2024; Larsson et al., 2024).
@@
for all $\alpha>0$, where the first inequality follows from Markov’s inequality and the second inequality from the definition of the e-value.
Recently, Koning (2024a) has shown that such a p-value is not just some conservative p-value, but has deep links to the data-dependent selection of the level and the expected relative discrepancy (see also Grünwald, 2024). In particular, an e-value-based p-value $p^{*}$ leads to decisions that have an expected discrepancy ratio below 1 for every data-dependent choice of the level $\alpha$. That is, for every possibly data-dependent level $\alpha$,
$\mathbb{E}\,r_{\alpha}=\mathbb{E}\left[\frac{\mathbb{P}(p^{*}\leq\alpha)}{\alpha}\right]\leq 1.$ (6)
A consequence is that a test $\phi_{\alpha}=\mathbb{I}\{p^{*}\leq\alpha\}$ based such a special p-value $p^{*}$ is valid for any data-dependent $\alpha$, in the sense that the expected relative discrepancy remains below 1. This is in stark contrast to traditional p-values, for which this error can be $\infty$, as discussed in Section 4. The connection is that a traditional $p$-value only satisfies (6) for every data-independent level.
While e-values are useful when $\alpha$ depends on the data, they should not be used when $\alpha$ is already fixed in advance. Indeed, in that case there is a considerable power loss, compared to using a traditional p-value.
@@ Discussion
The fundamental point of this paper is that significance testing works in practice because researchers all by default use the same $\alpha$. This ensures that $\alpha$ is independent of the data. If there are multiple standards for $\alpha$ in a field and a researcher chooses her $\alpha$ post hoc, then that spells the end of the validity of her tests. Her p-values may be perfectly valid and they will provide a pragmatic measure of the amount of proof against $H_{0}$ (Efron, 2010, p.31), but she will not have a test with the usual interpretation in terms of type I error control. Likewise, confidence intervals are no longer valid if $\alpha$ is chosen post hoc. As statisticians we should give this some careful thought. One solution would be that researchers fix $\alpha$ in advance and never change it, but they would not necessarily have enough incentives or awareness to do this.
@@
Existing papers that advocate decreasing $\alpha$ (Johnson et al., 2017; Szucs and Ioannidis, 2017; Benjamin et al., 2018; Ioannidis, 2018; Held, 2019; Schnog et al., 2021; Fitzpatrick et al., 2024), do not discuss the following question: would they already consider it an improvement if some people or journals in a field decrease $\alpha$, or do they consider it essential that the field as a whole uses the same $\alpha$? The present paper illustrates that these distinctions matter.
@@
Suppose that in a field the scenario occurs where some stick to $\alpha=0.05$ while others adopt $\alpha=0.005$ – thus providing an incentive for letting $\alpha$ depend on the data. One solution would be for journals to require pre-registration of the $\alpha$ before researchers conduct an experiment. Another partial solution would be to use e-values instead of p-values, as discussed in Section 5. Another way out is that a researcher uses a Bayesian approach, which does not require fixing an $\alpha$ and nevertheless tells us the probabilty that $H_{0}$ is false – which frequentist statistics never does (Kelter, 2020; Lakens et al., 2020; van de Schoot et al., 2021). A potential downside of Bayesian inference is that one needs to specify a prior distribution, which often depends on subjective considerations.
Appendix A Proofs
A.1 Proof of Proposition 1
Proof.
Since $\alpha=a_{2}\Leftrightarrow p>a_{1}$ and $p\leq a_{2}\Leftrightarrow\phi_{\alpha}=1$, under $H_{0}$ we have
$\mathbb{P}(\phi_{\alpha}=1|\alpha=a_{2})=$
$\mathbb{P}\big{\{}p\leq a_{2}|p>a_{1}\big{\}}=$
$\mathbb{P}\big{\{}a_{1}<p\leq a_{2}|p>a_{1}\big{\}}=$
$(a_{2}-a_{1})/(1-a_{1})=$
$a_{2}(1-a_{1}/a_{2})/(1-a_{1})<$
$a_{2}(1-a_{1})/(1-a_{1})=a_{2}.$
A.2 Proof of Proposition 2
Proof.
By (4), under $H_{0}$ we have
$r_{\alpha}=\mathbb{E}(\phi_{\alpha}/\alpha)=$
$\mathbb{P}(p\leq a_{1})\mathbb{E}(\phi_{\alpha}/\alpha|p\leq a_{1})+$
$\mathbb{P}(a_{1}<p\leq a_{2})\mathbb{E}(\phi_{\alpha}/\alpha|a_{1}<p\leq a_{2})=$
$a_{1}(1/a_{1})+(a_{2}-a_{1})(1/a_{2})=$
$1+(a_{2}-a_{1})/a_{2}.$
Appendix B Multiple hypotheses
Here we discuss scenarios where $\alpha$ is chosen post hoc and multiple hypotheses are tested. Consider the situation where a particular researcher tests multiple hypotheses $H_{1},...,H_{m}$ and computes corresponding p-values $p_{1},...,p_{m}$. To avoid a flood of false positive findings, it is common in this setting to apply a multiple testing corrrection. Here we will consider control of the familywise error rate (FWER) by use of the Bonferroni method.
The Bonferroni method takes the raw, i.e. unadjusted, p-values $p_{1},...,p_{m}$ as input and outputs adjusted p-values, which are simply $p_{1}^{b}=\min\{p_{1}\cdot m,1\},...,p_{m}^{b}=\min\{p_{m}\cdot m,1\}$. The method then rejects the hypotheses with adjusted p-values at most $\alpha$. Of course, an equivalent way to see which hypotheses are rejected is to simply check which of the raw p-values are at most $\alpha/m$.
The number of false positives, i.e. the number of true hypotheses that are rejected, if often denoted by $V$. The FWER is defined as $\mathbb{P}(V>0)$ and controlling the FWER means ensuring that $\mathbb{P}(V>0)\leq\alpha$. The Bonferroni method controls the FWER regardless of the dependence structure of the $m$ p-values and regardless of how many hypotheses are in fact false.
It should furter be remarked that the Bonferroni procedure can be made slightly more powerful through a stepwise procedure by Holm (Holm, 1979; Goeman and Solari, 2014). This method checks which hypotheses are rejected by Bonferroni and removes these from consideration. Then Bonferroni is applied to the hypotheses that are left, which can lead to additional rejections and this is repeated until a step is reached where there are no additional rejections. Here we will simply consider Bonferroni without the Holm improvement, since if Bonferroni does not control the FWER then by implication neither does Holm.
While there are many scenarios with data-dependent $\alpha$ where Bonferroni fails miserably, here we will focus on two scenarios, which bear slight resemblance to the scenarios considered in §2. As is usual, we will assume that the raw p-values corresponding to the true hypotheses are uniformly distributed on $[0,1]$. Further, for the sake of the below examples, we assume that $p_{1},...,p_{m}$ are independent of each other. Note that $FWER\leq\alpha$ is equivalent to $\mathbb{E}(\mathbbm{1}_{E})\leq\alpha$, where $E$ is defined to be the event that $V>0$. Further note that $\mathbb{E}(\mathbbm{1}_{E})\leq\alpha$ is equivalent to $\mathbb{E}(\mathbbm{1}_{E}/\alpha)\leq 1$, which we use below.
B.1 Example 1: two possible significance levels $\alpha$
As in the main paper, suppose the researcher chooses from two thresholds $10<a_{1}<a_{2}<$ 1. Let $p_{(1)}=\min\{p_{i}:1\leq i\leq m\}$ denote the smallest raw p-value and let $p_{(1)}^{b}=\min\{mp_{(1)},1\}$ be the smallest Bonferroni-adjusted p-value. Similar to before, one can image that a researcher is more likely to use the threshold $a_{1}$ if there is at least one rejection at $\alpha=a_{1}$, i.e. if $p_{(1)}^{b}\leq a_{1}$. To simplify, suppose she chooses $\alpha=a_{1}$ if $p_{(1)}^{b}\leq a_{1}$ and $\alpha=a_{2}$ otherwise, i.e. if $a_{1}<p_{(1)}^{b}\leq 1$. Like before, $\alpha$ depends on the p-value and thus the data, but we write $\alpha$ instead of $\alpha(X)$ for brevity.
B.1.1 No type I error control conditional on $\alpha$
Suppose all $m$ hypotheses are true and recall that $E=\{V>0\}$ is the event that there is at least one false positive finding. Then if $\alpha=a_{1}$, we have $p_{(1)}^{b}\leq a_{1}$. Thus $\mathbb{P}(E|\alpha=a_{1})$ is 1, i.e., the FWER is not controlled at all conditional on $\alpha=a_{1}$. The following proposition states that conditional on $\alpha=a_{2}$ it is the other way around, i.e. the test is conservative: $\mathbb{P}(E|\alpha=a_{2})<a_{2}$.
Proposition 3.
Suppose $m\geq 2$ and all hypotheses $H_{1},...,H_{m}$ are true. For the above procedure that takes $\alpha=a_{1}$ if $p\leq a_{1}$ and $\alpha=a_{2}$ otherwise, we have
$\mathbb{P}(E|\alpha=a_{2})=1-\Big{\{}(1-a_{2}/m)^{m}\big{/}(1-a_{1}/m)^{m}\Big{\}}<a_{2}.$
Proof.
Part 1. We have
$\mathbb{P}(E|\alpha=a_{2})=\mathbb{P}(p_{(1)}^{b}\leq a_{2}|p_{(1)}^{b}>a_{1})=\mathbb{P}(a_{1}<p_{(1)}^{b}\leq a_{2}|p_{(1)}^{b}>a_{1})=$
$\mathbb{P}(a_{1}<mp_{(1)}\leq a_{2}|mp_{(1)}>a_{1})=\mathbb{P}(a_{1}/m<p_{(1)}\leq a_{2}/m|p_{(1)}>a_{1}/m).$ (7)
If $m$ variables are independent and have a standard uniform distribution, the minimum of the $m$ variables has a $Beta(1,m)$ distribution (Hemerik et al., 2019), which has density function $f(x)=\{1/B(1,m)\}(1-x)^{m-1}$, where $B(1,m)$ is the normalization constant
$\int_{0}^{1}(1-x)^{m-1}dx=[-(1-x)^{m}/m]\Big{|}_{0}^{1}=0--1/m=1/m.$
Thus, $p_{(1)}$ has a $Beta(1,m)$ distribution, so that (7) equals
[…]
$\Big{\{}-(1-x)^{m}/m\Big{|}_{a_{1}/m}^{a_{2}/m}\Big{\}}\Big{/}\Big{\{}-(1-x)^{m}/m\Big{|}_{a_{1}/m}^{1}\Big{\}}=$
$\Big{\{}-(1-a_{2}/m)^{m}--(1-a_{1}/m)^{m}\Big{\}}\Big{/}\Big{\{}0--(1-a_{1}/m)^{m}\Big{\}}=$
$\Big{\{}(1-a_{1}/m)^{m}-(1-a_{2}/m)^{m}\Big{\}}\Big{/}\Big{\{}(1-a_{1}/m)^{m}\Big{\}}=$
$1-\Big{\{}(1-a_{2}/m)^{m}\big{/}(1-a_{1}/m)^{m}\Big{\}},$
as was to be shown.
Part 2. We now need to show that the above is smaller than $a_{2}$. This is equivalent to showing that
$(1-a_{2}/m)^{m}/(1-a_{1}/m)^{m}<1-a_{2},$
which is equivalent to
$(m-a_{2})^{m}(1-a_{2})^{-1}(m-a_{1})^{-m}<1.$
The left side is $1$ if $a_{2}=a_{1}$ and we need to show that it is strictly larger than 1 if $a_{2}>a_{1}$. We will do this by showing that the derivative of the function
$g(a_{2})=(m-a_{2})^{m}(1-a_{2})^{-1}(m-a_{1})^{-m}$
is strictly positive on $(0,1)$. We use the product rule and the chain rule to compute the derivative:
$\frac{d}{da_{2}}g(a_{2})=\Big{\{}(m-a_{2})^{m}(--(1-a_{2})^{-2})-m(m-a_{2})^{m-1}(1-a_{2})^{-1}\Big{\}}(m-a_{1})^{-m}.$
Thus, to show that $\frac{d}{da_{2}}g(a_{2})>0$, it suffices to show that
$(m-a_{2})^{m}(1-a_{2})^{-2}>m(m-a_{2})^{m-1}(1-a_{2})^{-1},$
i.e., that
$(m-a_{2})(1-a_{2})^{-1}>m.$
This is true since
$(1-a_{2}/m)(1-a_{2})^{-1}>1,$
so that we are done.
So far, we have seen that $\mathbb{P}(E|\alpha=a_{2})<a_{2}$ and $\mathbb{P}(E|\alpha=a_{1})=1>a_{1}$. Thus, conditional on $\alpha=a_{2}$ the method is conservative and conditional on $\alpha=a_{1}$ the method is much too liberal. Thus, the method does not provide FWER control conditional on the data-dependent $\alpha$.
B.1.2 No type I error control in terms of expected discrepancy
The above raises the question question whether we at least control the FWER on average. As in the main paper, we can consider the expected value of the discrepancy ratio, which in the context of the FWER is
$\mathbb{E}(r_{\alpha}):=\mathbb{E}\big{\{}\mathbb{E}(\mathbbm{1}_{E})/\alpha|\alpha\big{\}}=\mathbb{E}(\mathbbm{1}_{E}/\alpha).$
If this value is at most 1, then we control the FWER in an average sense. However, the following proposition implies that $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ is often larger than 1. The reason that it is not always larger than 1 is due to the fact that Bonferroni is conservative under independence if $\alpha$ is fixed, i.e., when we take $a_{1}=a_{2}$.
Proposition 4.
In the setting of Proposition 3, we have $\mathbb{E}(\mathbbm{1}_{E}/\alpha)=$
$\Big{\{}1-(1-a_{2}/m)^{m}\Big{\}}a_{2}^{-1}+\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}(a_{1}^{-1}-a_{2}^{-1}),$ (8)
which can be larger than $1$, depending on $m$, $a_{1}$ and $a_{2}$. The first term in (8) is smaller than 1 and represents $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ in the situation that $a_{1}$ and $a_{2}$ are identical, i.e. the case that $\alpha$ is fixed. The second term represents the increase of $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ due to $a_{2}$ being larger than $a_{1}$.
Proof.
We have
$\mathbb{E}(\mathbbm{1}_{E}/\alpha)=\mathbb{P}\Big{(}p_{(1)}^{b}\leq a_{1}\Big{)}(1/a_{1})+\mathbb{P}\Big{(}a_{1}\leq p_{(1)}^{b}\leq a_{2}\Big{)}(1/a_{2})=$
$\mathbb{P}\Big{(}p_{(1)}\leq a_{1}/m\Big{)}(1/a_{1})+\mathbb{P}\Big{(}a_{1}/m\leq p_{(1)}\leq a_{2}/m\Big{)}(1/a_{2})=$
Like in the proof of Proposition 3, $p_{(1)}$ has a $Beta(1,m)$ distribution, so that the above equals
[…]
Recall from the proof of Proposition 3 that $B(1,m)=1/m$. Consequently, the above equals
$ma_{1}^{-1}\big{\{}-(1-x)^{m}/m\big{\}}\Big{|}_{0}^{a_{1}/m}+ma_{2}^{-1}\big{\{}-(1-x)^{m}/m\big{\}}\Big{|}_{a_{1}/m}^{a_{2}/m}=$
$ma_{1}^{-1}\Big{\{}-(1-a_{1}/m)^{m}/m--(1-0)^{m}/m\Big{\}}+ma_{2}^{-1}\Big{\{}-(1-a_{2}/m)^{m}/m--(1-a_{1}/m)^{m}/m\Big{\}}=$
$a_{1}^{-1}\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}+a_{2}^{-1}\Big{\{}(1-a_{1}/m)^{m}-(1-a_{2}/m)^{m}\Big{\}}$
$a_{1}^{-1}\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}+a_{2}^{-1}\Big{\{}(1-(1-a_{2}/m)^{m})-(1-(1-a_{1}/m)^{m})\Big{\}}=$
$\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}(a_{1}^{-1}-a_{2}^{-1})+\Big{\{}1-(1-a_{2}/m)^{m}\Big{\}}a_{2}^{-1}.$ (9)
If we plug in $m=100$, $a_{1}=0.8$ and $a_{2}=0.81$, then the above is 0.70, rounded off to two decimals. This can be understood by noting that if we take $a_{1}=a_{2}=0.8$, we actually get a similar number, since we then consider standard Bonferroni with $\alpha=0.8$ and that method is particularly conservative under independence when $\alpha$ is large.
If we plug in $a_{1}=0.05$ and $a_{2}=0.1$, then (9) is 1.44, rounded off to two decimals. This illustrates that for many values of $a_{1}$ and $a_{2}$, (9) exceeds 1. ∎
For most reasonable settings, we find with Proposition 4 that $\mathbb{E}(\mathbbm{1}_{E}/\alpha)>1$. For example, if $m=1000$, $a_{1}=0.005$ and $a_{2}=0.05$ we find $\mathbb{E}(\mathbbm{1}_{E}/\alpha)\approx 1.87.$ Thus, usually Bonferroni will too liberal, in an average sense.
B.2 Example 2: infinitely many significance levels $\alpha$
Here, we again consider $0<C\leq 1$ and assume that if $p_{(1)}^{b}\leq C$, the researcher chooses $\alpha=p_{(1)}^{b}$. If $p_{(1)}^{b}>C$, she takes $\alpha=C$. Suppose that all $m$ hypotheses are true. Then, with this definition of $\alpha$, there is always at least one incorrect rejection if $p_{(1)}^{b}\leq C$. Thus, conditional on $\alpha<C$, $\mathbbm{1}_{E}$ is always equal to 1. Hence, we clearly do not have $\mathbb{E}(\mathbbm{1}_{E}/\alpha|\alpha<C)\leq 1$, i.e., we do not have FWER control conditional on $\alpha<C$.
Next we can investigate whether we at least control the FWER on average in the sense that $\mathbb{E}(r_{\alpha})=\mathbb{E}(\mathbbm{1}_{E}/\alpha)\leq 1.$ The following proposition states that this is not the case. In fact, $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ is infinity, like we saw in the case of one hypothesis.
Proposition 5.
Consider the setting of Propositions 3 and 4 above, but define $\alpha=p_{(1)}^{b}$ if $p_{(1)}^{b}\leq C$ and $\alpha=C$ otherwise. Then
$\mathbb{E}(r_{\alpha})=\mathbb{E}(\mathbbm{1}_{E}/\alpha)=\infty.$
Proof.
We will again use that $p_{(1)}$ has a $Beta(1,m)$ distribution, so has density function $f(x)=B(1,m)(1-x)^{m-1}$. Note that $E$ happens if and only if $p_{(1)}^{b}\leq C$, i.e., if and only if $p_{(1)}\leq C/m$. Hence, $\mathbb{E}(\mathbbm{1}_{E}/\alpha)=$
[…]
where we used that if $p_{(1)}^{b}\leq C$ then $1/\alpha=1/(mp_{(1)})$. Since $B(1,m)^{-1}=m$, the above equals
[…]
Note that $(1-x)^{m-1}\geq(1-C/m)^{m-1}$ for all $0\leq x\leq C/m$. Hence the above is at least
[…]

### v2

@@ 2.3 Example 2: Infinitely many significance levels $\alpha$
Now suppose there are not just two, but many thresholds to choose from. In fact, let us take this to the extreme by allowing the researcher to choose any threshold $0<\alpha\leq C$ up to some constant $C>0$, e.g. $C=0.05$ or $C=0.1$. If the researcher desires to report a discovery, it is easy to imagine that she is biased towards choosing an $\alpha$ that is at least as large as her p-value $p$, when possible. Again taking the most extreme choice, suppose she selects $\alpha=p$, unless $p>C$, in which case she selects $\alpha=C$.
@@
In case $\alpha<C$, $H_{0}$ is always rejected, so that we do not have conditional type I error rate control. Indeed, for every $a<C$, we have $r_{a}=1/a>1$. Next, we may ask whether we have type I error control not conditionally, but at least on average, in the sense that $\mathbb{E}r_{\alpha}\leq 1$ under $H_{0}$. Since $p$ is uniform on $[0,1]$ and $p\leq\alpha$ if and only if $p\leq C$, we have
$\mathbb{E}r_{\alpha}$ $=\mathbb{E}(\phi_{\alpha}/\alpha)$
$=\mathbb{P}(p>\alpha)\cdot 0+\mathbb{P}(p\leq\alpha)\mathbb{E}(\phi_{\alpha}/\alpha\mid p\leq\alpha)$
$=\mathbb{P}(p\leq\alpha)\mathbb{E}(\phi_{\alpha}/\alpha\mid\alpha=p\leq C)$
$=\mathbb{P}(p\leq\alpha)\mathbb{E}(1/p\mid p\leq C)$
@@
We see that the expected discrepancy ratio $\mathbb{E}r_{\alpha}$ is not only larger than 1, but in fact infinity. Thus, in expectation, the test is completely out of control.
@@ 3 Connection to e-values
In case we are only interested in $\mathbb{E}r_{\alpha}$, then a solution is offered by the e-value. An e-value is a measure of evidence that has been recently proposed as an alternative to the p-value [15, 42, 49, 11, 39, 27]. An e-value is typically defined as a non-negative random variable $e$, whose expectation is bounded by 1 under $H_{0}$:
$\mathbb{E}\,e\leq 1.$
In the context of a simple null and alternative hypothesis that only contain a single distribution, a prominent example of an $e$-value is the likelihood ratio between these distributions [42]. For a general null, desirable e-values are increasing functions of the likelihood ratio between the alternative and a kind of least-favorable distribution in the null [11, 31, 27].
@@
for all data-independent $\alpha>0$, where the first inequality follows from Markov’s inequality and the second inequality from the definition of the e-value.
Recently, [28] showed that $p^{*}$ is not just some overly conservative p-value, but exactly the kind of p-value that satisfies
$\mathbb{E}\,r_{\alpha}=\mathbb{E}\left(\frac{\mathbb{P}(p^{*}\leq\alpha\mid\alpha)}{\alpha}\right)\leq 1$
regardless of how $\alpha$ depends on the data. This is in stark contrast to traditional p-values, for which $\mathbb{E}r_{\alpha}$ can be $\infty$, as discussed in Section 2.3.
@@ Discussion
The fundamental point of this paper is that significance testing works in practice because researchers all by default use the same $\alpha$. This ensures that $\alpha$ is independent of the data [46]. If there are multiple standards for $\alpha$ in a field and a researcher chooses her $\alpha$ post hoc, then that spells the end of the validity of her tests. Likewise, confidence intervals are no longer valid if $\alpha$ is chosen post hoc. One solution would be that researchers fix $\alpha$ in advance and do not change it, but they would not necessarily have enough incentives or awareness to do this.
@@
Existing papers that advocate decreasing $\alpha$ [23, 43, 4, 22, 14, 41, 9], do not discuss the following question: would they already consider it an improvement if some people or journals in a field decrease $\alpha$, or do they consider it essential that the field as a whole uses the same $\alpha$? The present paper illustrates that these distinctions matter.
@@
Suppose that in a field the scenario occurs where some stick to $\alpha=0.05$ while others adopt $\alpha=0.005$—thus providing an incentive for letting $\alpha$ depend on the data. One solution would be to require pre-registration of $\alpha$ before researchers conduct an experiment [33]. Another partial solution would be to use e-values instead of p-values, as discussed in Section 3. Another way out is that a researcher uses a Bayesian approach, which does not require fixing an $\alpha$ and nevertheless tells us the probabilty that $H_{0}$ is false – which frequentist statistics never does [24, 30, 47]. A potential downside of Bayesian inference is that one needs to specify a prior distribution, which often depends on subjective considerations.